A student's grade in a course is determined by 4 quizzes and 1 exam. If the exam counts twice as much as each of the quizzes, what fraction of the final grade is determined by the exam?
Answer by ewatrrr(24785)

 
Hi,
x + x + x + x + 2x = 1
6x = 1
x = 1/6, fraction of the final grade determined by each quiz
Exam is 2(1/6) = 1/3
1/6 + 1/6 + 1/6 + 1/6 + 1/3 = 4/6 + 1/3 = 1